OpenStudy (anonymous):

Algerbra 1 help please! :) 28. Tell whether the sequence is arithmetic. If it is, what is the common difference? –12, –7, –2, 3, . . . (1 point) yes; 5 yes; 8 yes; 3 no

OpenStudy (anonymous):

it suppose to be yes,5 yes,8 yes,3 no

OpenStudy (tkhunny):

Subtract successive terms: –12, –7, –2, 3 -7 - (-12) = -7 + 12 = 5 -2 - (-7) = -2 + 7 = 5 3 - (-2) = 3 + 2 = 5 Assuming that this pattern continues, we can answer the question.

OpenStudy (anonymous):

I dont really know what arithmetic means... I'm guessing yes,5. @tkhunny

OpenStudy (tkhunny):

Seriously? I always wonder why problems are given without ANY information that might allow the student to succeed. Arithmetic Sequence Arithmetic Progression An ordered set of numbers such that the difference of any two consecutive numbers is constant. In other words, there is a common difference in advancing from one value to the next. The difference is always the same.

OpenStudy (anonymous):

@tkhunny oh ok thanks so much! I get it now.... :D